OpenStudy (angela210793):

I've been studying geometry from 11am-10pm(now) and the result I got isn't the one it should be....Please help me!!! PLEASE ^_^ In the triangle ABC AB=8 and BC=12,the angle between AB and BC is 60 degrees.The triangle rotates around BC.Find the volume of the 'thing' formed from this rotation :)

OpenStudy (amistre64):

rotates around bc? really

OpenStudy (angela210793):

O.o yes it does...I can draw a pic if u want....

OpenStudy (anonymous):

k here i will tell u how to do it when u rotate a triangle about 1 of its sides u get a figure with two cones of same radius attached now try to find the hight of each cone and u have the radius which is hight of the triange from side BC so u will have the volume by the formula V of cone = (1/3)(pi)(r^2)h

OpenStudy (amistre64):

like in a circle?

OpenStudy (angela210793):

yea i know that thnx :) but the problem is that i am not sure...we can't count the area of the circle formed twice can we?

OpenStudy (amistre64):

its not gonna count twice

OpenStudy (angela210793):

so it must be V=V1+V2-piR^2?

OpenStudy (angela210793):

or not?

OpenStudy (amistre64):

volume top cone + volume bottom cone = volume of rotation

OpenStudy (anonymous):

amistre is correct

OpenStudy (amistre64):

when you press your hands together; do you subtract the volume created by them touching?

OpenStudy (amistre64):

you add the volume of one to the other;

OpenStudy (angela210793):

how abt the circle? O.o

OpenStudy (anonymous):

a circle has a area not a volume

OpenStudy (amistre64):

the circle is just a way to measure the volume for the top and the volume for the bottom

OpenStudy (angela210793):

yea but it will be counted twice since the same cones have the same base right?

OpenStudy (amistre64):

when....bases..touch.... they dont count as ....twice anything

OpenStudy (amistre64):

when you press your hands together; no EXTRA volume is created to take away

OpenStudy (amistre64):

when you stack a can of corn on top of another can of corn; you havent add any EXTRA volume by staking

OpenStudy (amistre64):

the place where they touch, is just that .... the place where they touch

OpenStudy (amistre64):

prolly what your thinking of is that you are used to seeing area as a surface that has some depth to it; when area is infinitely thin has zero depth

OpenStudy (anonymous):

How the hell do you form 2 cones by rotating about one of the edges? The tip of the triangle (point A) will make a circle through space...

OpenStudy (amistre64):

an edge make an axel in space

OpenStudy (amistre64):

ever play with those claker balls on a stick?

OpenStudy (anonymous):

Right, so all points on edges AB, an AC will form circles when rotated all the way about the BC axis

OpenStudy (amistre64):

right

OpenStudy (anonymous):

oooh, ok i was thinking about it wrong, i was thinking 2 cones with their tips being point A

OpenStudy (anonymous):

why is there 2 cones?

OpenStudy (anonymous):

ABC, AB = 8, BC = 12 < ABC = 60

OpenStudy (amistre64):

becasue the triangle is spun on an edge makind a top and bottom cone

OpenStudy (anonymous):

The circle that A makes through space is the base of both cones, One cone is formed by each half of the triangle rotating 360 degrees through space

OpenStudy (amistre64):

i get a radius of 7; and a height of 4 for the top i get a radius of 7 and a height of 8 for the bottom part

OpenStudy (anonymous):

you mean a right and left cone, not top and bottom

OpenStudy (amistre64):

does it really matter how we hold the edge?

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

(1/3) pi 49(4) + (1/3) pi 49(8) should be the volume round abt

OpenStudy (amistre64):

615.75.... is the best I can do with truncating numbers

OpenStudy (anonymous):

i dont follow the rotation, also your triangle doesnt make sense, can you label things

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

ok i drew my triangle a little differently , so first maybe its smarter to actually find the other angles and sides, so we dont just guess the shape

OpenStudy (amistre64):

mines not to scale :)

OpenStudy (anonymous):

Think about a flag, that is a right triangle. rotate that flag around the flag pole. You get one cone. The triangle we are dealing with is not a right triangle but can be divided in half and considered as 2 right triangles. So 2 cones.

OpenStudy (anonymous):

so i get 8 sin 60 for the radius , 8 sin pi/3

OpenStudy (anonymous):

8 sin pi/3 ~ 6.92

OpenStudy (amistre64):

if we want accuracy; it should go \[\frac{pi (8sin(30))^2*4}{3}+\frac{pi (8sin(30))^2*8}{3}\]

OpenStudy (anonymous):

how do you get 4

OpenStudy (anonymous):

oh 8 sin pi/6

OpenStudy (amistre64):

lol.... i did sin instead of cos :)

OpenStudy (amistre64):

8sin(30) = height of top = 4

OpenStudy (amistre64):

12 - 4 = 8 for height of bottom

OpenStudy (anonymous):

ok , so change sin to cos in your formula

OpenStudy (amistre64):

yes ...

OpenStudy (anonymous):

ok looks good !

OpenStudy (amistre64):

we can clean it up by factoring: \[\frac{4\pi (8cos(30))^2}{3}(1+2)\] right?

OpenStudy (anonymous):

16 /3 pi ( 8 cos (pi/6))^2

OpenStudy (anonymous):

woops

OpenStudy (amistre64):

\[4\pi\ (8 cos(30))^2\] seems to be the simplest form

OpenStudy (anonymous):

4/3 pi ( 8 cos ( pi/6))^2

OpenStudy (anonymous):

grrr

OpenStudy (amistre64):

:) 603.185.....

OpenStudy (anonymous):

this interface is not forgiving

OpenStudy (anonymous):

can i ask a quick question, given a focus and directrix, there is only one e such that e = 1 , so there is only one parabola

OpenStudy (amistre64):

we dint use eccentricity much, so Id have to look it up; but libraries closing so I gotta move to the veranda :)

OpenStudy (anonymous):

but there are many ellipses and parabolas

OpenStudy (anonymous):

ok bbl ?

OpenStudy (angela210793):

:(:(:(:(I'm sooooo sorry guys..:(:( chrome wasn't responding :(:(I'm sorry :(

OpenStudy (angela210793):

Hmmmm......the V=192pi tht's wht the answer must be acc.to my math book

OpenStudy (angela210793):

Thanks for trying anyway :) ^_^